# Multi-modal Gaussian Process Variational Autoencoders for Neural and Behavioral Data

### v1

@@ 6 Conclusion
Pre-processing: Original synthetic visual stimuli were sampled at 125 Hz and the neural and torque recordings were sampled at 10K Hz [40, 41, 42].
To prepare these data for evaluation with the MM-GPVAE, we first downsampled neural information and torque measurements to 1K Hz by binning the spike counts and averaging the torque measurements at this temporal resolution. To align these measurements with the visual stimuli, we upsampled the images 8-fold. The entire recording was 20 seconds long and was split into 6 evenly-divided trials.
The visual stimuli contained 3024 pixels and there were ten total hawkmoth motor neurons. We set the neural-independent subspace to 2-dimensional, the images-independent subspace to 1-dimensional, and an additional 1 dimension for the shared subspace. To encourage slow-evolving smooth latents in the shared and image subspaces, and faster-evolving neural latents, we initialized the length scale parameters for each latent dimension to different values. The length scale was set to 10 for the neural latents, 150 for the shared latent, and 300 for the image latent. This biased the model fits to capture slower dynamics in the image subspace and the faster dynamics in the neural subspace.
@@
We found that how we initially set the value of could effect the performance of the MM-GPVAE, especially regarding the reconstruction of the behavioral data. Although we learned the value jointly with the other parameters during optimization, the value of tended to vary minimally from its initial value. Therefore, our setting of tended to act as a reconstruction penalty, balancing the ability of the MM-GPVAE to prefer to reconstruct either the behavioral or neural data. Such a scaling term in the ELBO reconstruction has been seen in other models [11], and in the original GP-VAE this parameter was chosen carefully through a specific cross-validation approach [24]. Here, we simply set to a value proportional to the dimensionality of the behavioral modality, which tended to balance the neural and behavioral terms in the ELBO, and generate good reconstructions for each modality. The inital values for were 1000, 100, 1e-6, and 1 for GP-VAE (simulated), MM-GPVAE (simulated), MM-GPVAE (fly), and MM-GPVAE (moth) respectively.
@@
Here we show the derivation of the evidence lower bound used for the MM-GPVAE. For clarity, we will start by deriving the ELBO for the MM-GPVAE just in the time domain. We set the GP prior parameters , to be the loadings matrix and offsets from equation 4 in the manuscript, to be the observed timepoints, and to be the collection of all the shared and independent latents .

### v2

@@ 7 Ethics Statement
Pre-processing: For our hawkmoth data, the original synthetic visual stimuli were sampled at 125 Hz and the neural and torque recordings were sampled at 10K Hz (Sprayberry and Daniel, 2007; Sikandar et al., 2023; Putney et al., 2019). To prepare these data for evaluation with the MM-GPVAE, we first downsampled neural information and torque measurements to 1K Hz by binning the spike counts and averaging the torque measurements at this temporal resolution. To align these measurements with the visual stimuli, we upsampled the images 8-fold. The entire recording was 20 seconds long and was split into 6 evenly-divided trials. Here, it is worth noting that the trials are much longer ( 3300 timebins) than in the other experiments in this work.
The visual stimuli contained 3024 pixels and there were ten total hawkmoth motor neurons. We set the neural-independent subspace to 1-dimensional, the images-independent subspace to 1-dimensional, and an additional 1 dimension for the shared subspace. To encourage slow-evolving smooth latents in the shared and image subspaces, and faster-evolving neural latents, we initialized the length scale parameters for each latent dimension to different values. The length scale was set to 10 for the neural latents, 150 for the shared latent, and 300 for the image latent. This biased the model fits to capture slower dynamics in the image subspace and the faster dynamics in the neural subspace.
@@
To emphasize the added interpretability we get from the linear layer in the neural likelihood and the role of the shared latent in generating the data from each modality, in Figure 21 we show the reconstructions for both modalities using one or no shared latent variables (Fig 21 (a)) as well as 2 example neurons’ tuning to the neural and shared latent. Here we can see that neuron 5 is more positively modulated by the shared latent while neuron 9 is negatively modulated by the shared latent.
@@
Latent dimensionality selection: In the case of the simulated examples in this work, the dimensionality was always set to the true generative dimensionality. For the real neural dataset examples in the main paper, the dimensionality was chosen using a distinct approach for each modality. For the neural dataset, we use PCA and select the number of dimensions that capture greater than 97% of the variance. For the other modality (limb position or visual image) the dimensionality was chosen by increasing the number of dimensions in a unimodal Fourier GP-VAE until cross-validated performance did not show significant improvement. Then, we identify the appropriate number of shared dimensions by fitting the full MM-GPVAE model to both modalities and systematically increasing the number of dimensions shared between datasets until predictive performance no longer improved. Across both datasets, we always saw that adding some shared dimensions increased overall predictive performance compared to none, suggesting shared structure across the data modalities for these experiments.
We also note that the dimensionality selection procedure was used to approximate shared and independent dimensionality as best as possible, but across CV folds the variance in cross validated log-likelihood was high. However, small changes to the number of shared and independent dimensions did not change latent representations or decoding performance much in our models.
For the evaluation where we compare MM-GPVAE to TNDM and PSID, we choose approximately the same dimensionality as in the original TNDM paper (Hurwitz et al. (2021)) so readers can easily compare fits across models.
@@
We found that how we initially set the value of could effect the performance of the MM-GPVAE, especially regarding the reconstruction of the behavioral data. Although we learned the value jointly with the other parameters during optimization, the value of tended to vary minimally from its initial value. Therefore, our setting of tended to act as a reconstruction penalty, balancing the ability of the MM-GPVAE to prefer to reconstruct either the behavioral or neural data. Such a scaling term in the ELBO reconstruction has been seen in other models (Hurwitz et al., 2021), and in the original GP-VAE this parameter was chosen carefully through a specific cross-validation approach (Casale et al., 2018). Here, we simply set to a value proportional to the dimensionality of the behavioral modality, which tended to balance the neural and behavioral terms in the ELBO, and generate good reconstructions for each modality. The initial values for were 1000, 100, 1e-6, and 1 for GP-VAE (simulated), MM-GPVAE (simulated), MM-GPVAE (fly), and MM-GPVAE (moth) respectively.
@@
Here we show the derivation of the evidence lower bound used for the MM-GPVAE. For clarity, we will start by deriving the ELBO for the MM-GPVAE in the time domain, and ignore the Fourier representation in this derivation. The derivation trivially applies to the Fourier representation as well, as the mapping between the two domains is linear. We set the GP prior parameters , to be the loadings matrix and offsets from equation 4 in the manuscript, to be the observed timepoints, and to be the collection of all the shared and independent latents .